OpenStudy (anonymous):

Use the discriminant to determine the number and type of solutions the equation has. x2 + 8x + 10 = 0 A. no real solutions B. one real solution C. two rational solutions D. two irrational solutions

OpenStudy (zepp):

Okay, first, what's a discriminant?

OpenStudy (zepp):

In the quadratic formula, the discriminant is \(\sqrt{b^2-4ac}\)

OpenStudy (zepp):

Here, \(x^2 + 8x + 10 = 0\) a=1 b=8 c=10 Replace a,b,c in the discriminant by those numbers and tell me what you'll get.

OpenStudy (anonymous):

@jim_thompson5910 , may you help me ?

jimthompson5910 (jim_thompson5910):

did zepp's explanation make sense?

OpenStudy (anonymous):

Not really .

OpenStudy (anonymous):

I think its B

jimthompson5910 (jim_thompson5910):

The discriminant is D = b^2-4ac

OpenStudy (anonymous):

Oh . The answer is D ? Thats what your saying ?

jimthompson5910 (jim_thompson5910):

In the case of x^2 + 8x + 10, we see that a = 1, b = 8 and c = 10 So D = b^2-4ac D = (8)^2-4(1)(10) D = 24

jimthompson5910 (jim_thompson5910):

If the discriminant is positive, what does that mean?

OpenStudy (anonymous):

Im not sure

jimthompson5910 (jim_thompson5910):

If the discriminant is positive, then you'll have two real solutions. Since the discriminant is NOT a perfect square, this means that you'll have 2 irrational solutions.

OpenStudy (anonymous):

OH . So D !!

jimthompson5910 (jim_thompson5910):

yes you got it

OpenStudy (zepp):

Um, so how my explanation doesn't make sense? :(